OpenStudy (anonymous):

x^2 =(x+y)÷(x−y) find dy/dx in two ways: (i) by multiplying (x-y) on both sides (ii) by quotient rule Do you get the same result each way? please give me explanation if it isn't same and how to verify if both ways is the same.

OpenStudy (anonymous):

yes you do. but it is a pain to make it so

OpenStudy (anonymous):

what kind of a question is "do you get the same result"???!!!

OpenStudy (anonymous):

by doing (i) and (ii). do you get the same answer?

OpenStudy (anonymous):

first of all i assume you are looking for \[y'\] yes?

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

\[x^2=\frac{x+y}{x-y}\] \[2x=\frac{(x-y)(1+y')-(x+y)(1-y')}{(x-y)^2}\]

OpenStudy (anonymous):

gotta solve this for \[y'\] yikes. i need pencil and paper

OpenStudy (anonymous):

you probably dont get the same result both times

OpenStudy (anonymous):

because the first way of writing it has a denominator, so the points where denominator equals zero should be excluded

OpenStudy (anonymous):

the second expression doesnt have a division.

OpenStudy (anonymous):

well this is what i got the first time @elecengineer please check: \[2x(x-y)^2=2xy'-2yy'\] \[2x(x-y)^2=y'(2x-2y)\] \[y'=\frac{2x(x-y)^2}{2x-2y}=\frac{x(x-y)^2}{x-y}=x(x-y)\]

OpenStudy (anonymous):

after a bunch o'algebra

OpenStudy (anonymous):

i don't have to worry about division by 0 because in the original expression you cannot have \[x=y\]

OpenStudy (anonymous):

let me try second way.

OpenStudy (anonymous):

yup.. ==" calm down.. look slowly..

OpenStudy (anonymous):

@elecengineer it is the first version that has not denominator. second one does. now i have to figure out if i made a mistake, or why they are the same. can i cancel?

OpenStudy (anonymous):

oooooooooooooooooops

OpenStudy (anonymous):

\[x^3-x^2y=x+y\] maybe this works better

OpenStudy (anonymous):

\[3x^2 -2xy-x^2y'=1+y'\] \[3x^2-2xy-1=y'+x^2y'=y'(1+x^2)\] \[y'=\frac{3x^2-2xy-1}{1+x^2}\]

OpenStudy (anonymous):

better algebra this time?

OpenStudy (anonymous):

Digesting it.......

OpenStudy (anonymous):

well what is going to be really annoying is figuring out why they are the same.

OpenStudy (anonymous):

oh lord i think i made a mistake in the first one. let me try it again

OpenStudy (anonymous):

i thought it is annoyting to think that why the answer are not the same?

OpenStudy (anonymous):

let me check the first solution

OpenStudy (anonymous):

okay.

OpenStudy (anonymous):

ok i messed up and wrote something wrong on my paper.

OpenStudy (anonymous):

yup.. i think if i wasn't mistaken.. the first equation you had wrong druring the 3rd line..

OpenStudy (anonymous):

we have \[x^2=\frac{x+y}{x-y}\] \[2x=\frac{(x-y)(1+y')-(x+y)(1-y')}{(x-y)^2}\] \[2x(x-y)^2=(x-y)(1+y')-(x+y)(1-y')\]

OpenStudy (anonymous):

somehow i screwed up expanding on the right hand side. it is \[x+xy'-y-yy'-(x-xy'+y-yy')\] \[x+xy'-y-yy'-x+xy'-y+yy'\] \[=2xy'-2y\]

OpenStudy (anonymous):

yup! i got that too!

OpenStudy (anonymous):

so now we have \[2x(x-y)^2=2xy'-2y\] \[2x(x-y)^2+2y=2xy'\] \[\frac{2x(x-y)^2+2y}{2x}=y'\] \[y'=\frac{x(x-y)^2+y}{x}\] unless i made another mistake

OpenStudy (anonymous):

i think you got it right.. but now i was wondering.. why isn't the answer is same.. LOL

OpenStudy (anonymous):

and i have no idea why they are the same. oh they are the same

OpenStudy (anonymous):

you just can't see it. don't forget that you have an equality to begin with . it will depending on using it at some point

OpenStudy (anonymous):

you started with \[x^2=\frac{x+y}{x-y}\] so somehow using that you can show these are the same

OpenStudy (anonymous):

Ermmmmmm.. digesting...

OpenStudy (anonymous):

i guess what i am saying is that if you look in one expression for the derivative you see \[x^2\] so you are free to replace is by \[\frac{x+y}{x-y}\] if you choose. there is probably some clever algebra to be done, but i don't see it

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

yw. sorry about the algebra. if i had an hour i bet i could do it....maybe

OpenStudy (anonymous):

hahaha.. its okay.. you already gave me a big help.

OpenStudy (anonymous):

Can you do me 1 more favour? Can you show me the way to verify both (i) & (ii) are the same?

OpenStudy (anonymous):

@satellite73 hlp!!><

OpenStudy (amistre64):

x^2 =(x+y)/(x−y) find dy/dx in two ways: (i) by multiplying (x-y) on both sides x^2 (x-y) = x+y Dx(x^3-x^2 y = x+y) 3x^2 -x^2 y' -2xy = 1+y' 3x^2 -2xy -1 = y' +x^2 y' 3x^2 -2xy -1 = y'(1 +x^2) y' = (3x^2 -2xy -1)/(1+x^2) (ii) by quotient rule x^2 =(x+y)/(x−y) Dx [x^2 =(x+y)/(x−y)] 2x = [(x-y)(1+y') - (x+y)(1-y')]/[x-y]^2 2x(x-y)^2 = x+xy' -y -yy' -x +xy' -y +yy' 2x(x-y)^2 = xy' -2y +xy' 2x(x-y)^2 +2y = y'(2x) y' = (x-y)^2 + y/x ...................................... (3x^2 -2xy -1)/(1+x^2) = (x-y)^2 +yx ; is this what we got so far?

OpenStudy (anonymous):

yaya!

OpenStudy (amistre64):

well, they dont make the same graph when plotted ....

OpenStudy (anonymous):

and then... =.=" stucked... this question is wrong??

OpenStudy (anonymous):

nothing went wrong

OpenStudy (anonymous):

they are surely the same. but to do that you are going to have to do a raft of algebra. i bet amistre can do it yes?

OpenStudy (anonymous):

Totally confused... o.O"

OpenStudy (amistre64):

i cant see any errors .. but i cant see the equality either :)

OpenStudy (anonymous):

no no you did nothing wrong

OpenStudy (anonymous):

i made a couple errors but fixed them and i got what you got

OpenStudy (amistre64):

not without buying a few notebooks maybe :)

OpenStudy (anonymous):

this doesn't help me.

OpenStudy (anonymous):

we have two expressions for y ' yes?

OpenStudy (amistre64):

yes ....

OpenStudy (amistre64):

I even tried plugging in the same x and y value into each of them .... to test

OpenStudy (anonymous):

\[y'=\frac{3x^2 -2xy -1}{1+x^2}\] and \[y'=(x-y)^2+\frac{y}{x}=\frac{x(x-y)^2}{x}\]

OpenStudy (anonymous):

now i am next to positive they are the same. not that they are the same expression, because they certainly are not

OpenStudy (anonymous):

but don't forget we started with \[x^2=\frac{x+y}{x-y}\] so using this identity and some algebra that i don't know, i am willing to bet you can show these are the same

OpenStudy (amistre64):

\[y'=(x-y)^2+\frac{y}{x}=\frac{x(x-y)^2+y}{x}\] right?

OpenStudy (anonymous):

that is what i got yes

OpenStudy (amistre64):

oh good; then you musta dropped a y in transit :)

OpenStudy (anonymous):

oh sorry yes a "typo"

OpenStudy (anonymous):

so for example in the expression \[1+x^2\] in the denominator of first example, you can replace \[x^2\] by \[\frac{x+y}{x-y}\] if you like

OpenStudy (anonymous):

let me try to do some algebra. it will be a pain for sure but i cannot do it here. somehow you can show that these are the same, given the original identity

OpenStudy (anonymous):

Brain storm had turn to be a storm in the brain... =."

OpenStudy (anonymous):

i must be really messing up because i got the derivative is 1.

OpenStudy (anonymous):

\[y'=(x-y)^2+\frac{y}{x}\]

OpenStudy (anonymous):

everyone should give you medals, your first post was 2 hours ago o.O

OpenStudy (anonymous):

and \[x^2=\frac{x+y}{x-y}\rightarrow x^2(x-y)=x+y \rightarrow x-y=\frac{x+y}{x^2} \rightarrow \] \[(x-y)^2=\frac{(x+y)(x-y)}{x^2}=\frac{x^2-y^2}{x^2}\] \[y'=\frac{x^2-y^2}{x^2}+\frac{y}{x}=\frac{x^2-y^2+xy}{x}\]

OpenStudy (anonymous):

ok joemath i have to go. your assignment, if you decide to accept it, so to show that these derivatives are the same: \[y'=\frac{x(x-y)^2+y}{x}\] and \[\frac{3x^2-2xy-1}{x^2+1}\]

OpenStudy (anonymous):

you are going to need that \[x^2=\frac{x+y}{x-y}\]

OpenStudy (anonymous):

good luck. as always, should you or any of your IM force be caught or killed, the sensei will disavow any knowledge of your actions.

OpenStudy (anonymous):

if you do this in two minutes i am never coming back to this site.

OpenStudy (anonymous):

lolol, im at school right now, i'll work on this during my american lit class

OpenStudy (anonymous):

have fun

OpenStudy (anonymous):

i got the same thing as you using the quotient rule

OpenStudy (zarkon):

solve for y in the equation \[x^2=\frac{x+y}{x-y}\] getting \[y=\frac{x(x^2-1)}{x^2+1}\] now substitute this into both derivatives and simplify...you will see that they are the same

myininaya (myininaya):

read and weep